OpenStudy (anonymous):

Amanda teaches the art of quilling to 4 students. These students each teach the art of quilling to 4 other students. If this process continues for 5 generations, people will know the art of quilling.

OpenStudy (anonymous):

In each generation the number of people who know the art of quilling is multiplied by 4. gen 1 = 1 gen 2 - 1 x 4 = 4 gen 3 - 4 x 4 = 16 gen 4 - 16 x 4 = 6 gen 5 - 64 x 4 = 256 So after five generations 256 people should know the art of quilling. Unless generation 1 doesn't count and then it would be 1024 people.

OpenStudy (anonymous):

gen one doesnt count. I got it though thank you